Light-emitting diode (LED) light bulbs have become required in recent years, but do they make financial sense? Suppose a typical 60-watt incandescent light bulb costs $.45 and lasts for 1,000 hours. A 7-watt LED, which provides the same light, costs $2.25 and lasts for 40,000 hours. A kilowatt-hour of electricity costs $.121, which is about the national average. A kilowatt-hour is 1,000 watts for 1 hour. However, electricity costs actually vary quite a bit depending on location and user type (you can get information on your rates from your local power company). An industrial user in West Virginia might pay $.04 per kilowatt-hour whereas a residential user in Hawaii might pay $.25.

You require a 10 percent return and use a light fixture 500 hours per year. What is the break-even cost per kilowatt-hour?
Answer #1

For Break Even cost set the EAC of both light bulbs equal to solve for cost.

EAC = NPV/Annuity factor

NPV= - (Initial Price + Cost per year at annuity factor)

Cost per year = Kilowatts per hour * total hours used in a year * cost per kilowatt

Incandescent bulb t= 1000/500 = 2

LED= 40000/500 = 80

Kilowatts per hour = W/1000

EAC of 60-watt incandescent light bulb= - (Initial Price + Cost per year at annuity factor)/ annuity factor

[- 0.45 - (60/1000*500*C)(PVIFA10%,2)]/(PVIFA10%,2) = [-2.25- (7/1000*500*C)(PVIFA10%,80)] / (PVIFA10%,80)

[- 0.45 - ((60/1000)*500*C)(1.7355)]/(1.7355) = [-2.25- ((7/1000)*500*C)(9.9951)] / (9.9951)

Solving for C, break even cost per KW hour = $0.0068